OpenStudy (anonymous):

A largest pumped storage power operate allowing water to fall to a lower reservoir through a drop of 500m, passing through six turbines on the way which generate 310 MW of power. Q.water has a density of 1000kg/m^3, what is the minimum volume of water that must flow through each turbine each second?

OpenStudy (stormfire1):

First calculate the mass of water required to generate 310 MJ: \[3.1x10^8=(9.8m/s^2)(500m)mass\]\[mass=6.327x10^4kg\]Now that you have the mass calculate the volume:\[\frac{6.327x10^4kg}{1000kg/m^3}=6,327m^3\]

OpenStudy (anonymous):

We are asked about volume of water flowing through each turbine, so we must divide the last number by 6, cause there are 6 turbines and they all generate this power together.

OpenStudy (stormfire1):

@ArchiePhysics: Good catch...I missed the "each" part in the last sentence. @bebep207: The final answer above needs to be divided by 6 which leaves you with: \[1054.5m^3 ~of~water~per~turbine\]

OpenStudy (anonymous):

stormfire1, ure the best x). ty archie for the correction